OpenStudy (anonymous):

graph a line with the indicated slope and y-intercept. Slope=-1 y-intercept (0,10)

OpenStudy (anonymous):

Need help

OpenStudy (anonymous):

y=mx+c slope intercept form c is y-intercept =10 y=(-1)x+10 y=-x+10 now put the values of x and calculate y draw graph have you got it